OpenStudy (anonymous):

what is Tan of -90? :/

OpenStudy (anonymous):

\[\tan x=\frac{ \sin x }{ \cos x }\] Now at -90 or +90 cosx=0 And we cannot divide by 0, this means that we do not know exactly what happens at 0.

OpenStudy (anonymous):

tan(-90)=-tan90=-infinite as tan90=sin90/cos90 and cos90=0 also 1/0=infinte so we get -infinite